the intensity level of two sounds are 100dB and 50dB, ?hat is the ratio of their intensities?


1
Expert's answer
2021-07-28T03:29:04-0400

The intensity level of the sound

"\beta=10\lg\frac{I}{I_0}"

Hence

"\beta_2-\beta_1=10\lg\frac{I_2}{I_1}"

"\frac{I_2}{I_1}=10^{(\beta_2-\beta_1)/10}=10^{(100-50)/10}=100000"